The score 28 and 40 bound the middle 95% of the distribution. What percent is left at the tails of the distribution


1
Expert's answer
2022-06-01T11:02:32-0400
"100\%-95\%=5\%"

5% is left at the tails of the distribution.


"P(Z<z_1)=0.025"

"z_1=-1.96"


"P(Z>z_2)=0.025"

"z_1=1.96"

"\dfrac{28-\mu}{\sigma}=-1.96"

"\dfrac{40-\mu}{\sigma}=1.96"

Then


"\mu-28=40-\mu"

"\mu=34"

"\dfrac{40-34}{\sigma}=1.96"

"\sigma=\dfrac{6}{1.96}=3.0612"